OpenStudy (anonymous):

Solve the system using any method: 4x - 3y = -13 x + 3y = -7 A. x = -4, y =- 1 B. x = -4, y = 4 C. x = -1, y = 4 D. x = 4, y = -1

OpenStudy (anonymous):

@biggirl99 try substituting each one and see which makes the equations true

OpenStudy (anonymous):

4(-4)-3(-1)=13. true or false?

OpenStudy (anonymous):

sorry, -13, not 13

OpenStudy (anonymous):

false

OpenStudy (anonymous):

what does it equal, then?

OpenStudy (anonymous):

I did it in my calculator and it said 12

OpenStudy (anonymous):

oh okay sorry I must have put it in wrong or something

OpenStudy (anonymous):

true

OpenStudy (anonymous):

wait, ignore. wrong one

OpenStudy (anonymous):

-4+3(-1)=-7 true or false?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

3x-1=-3. -4+(-3)=-7

OpenStudy (anonymous):

sorry, i gtg. but i think you got it. since both of those are true, the answer is A.

OpenStudy (anonymous):